OpenStudy (anonymous):

Two Martians, Splott and Fizzle, have solved the equation 2x + 4 = –3x + 14. Examine the work of Splott and Fizzle. Identify any errors in the Martians’ calculations and explain, using complete sentences, what corrections they should make. Splott Fizzle 2x + 4 = –3x + 14 2x + 4 = –3x + 14 2x + 4 + 3x = –3x + 14 + 3x 2x + 4 – 2x = –3x + 14 – 2x 5x + 4 = 14 4 = 14 – 5x 5x + 4 + 4 = 14 + 4 4 – 14 = 14 – 5x – 14 5x = 18 –12 = –5x 5 x divided by 5 is equal to 18 divided by 5 –12 + 5 = –5x + 5 x = 18 over 5 –7 = x

OpenStudy (igreen):

I will solve it, and you tell me what mistakes(if they made any) they made.

OpenStudy (igreen):

2x + 4 = –3x + 14 Subtract 4 to both sides: 2x = -3x + 14 - 4 Subtract: 2x = -3x + 10 Add 3x to both sides: 2x + 3x = 10 Add: 5x = 10 Divide 5 to both sides: x = 10 / 5 Divide: x = 2

OpenStudy (igreen):

@taylor12344

OpenStudy (igreen):

You should take a Print Screen of the question so I can see it easier.

OpenStudy (igreen):

@taylor12344